bread is being produced being produced in two bakeries in lusaka. they compete keenly and there is no collusion between them -explicit or implicit / the bread they produce is identical
X1 represents the output of the first bakery
X2 represents the output of the second bakery
the marginal cost in ngwee of the firm is described by
MC1 =400+1/2X
MC2 =2 X2+100

the demand for bread in lusaka is 4500 dozen loaves a day (at any price)
(a) what equation is needed to complete the model?
(b) how many loaves of bread will be produced by bakery 1 and 2
(c) at what price(s) will bread be sold in the two bakeries?
1
Expert's answer
2019-07-31T11:02:13-0400

Output of the first bakery, q1= X1

Output of the second bakery, q2= X2

MC1 =400+1/2X

MC2 =2 X2+100 

Demand, p1= 4500

Demand, p2= 4500

a. MC(q) = MR(q), but for PD monopolist MR(q) = p(q). PD monopolist produces at perfectly competitive outcome where p(q) = MC(q)

b. How many loaves of bread will be produced by bakery 1 and 2

Bakery 1= 1x= 400+ o.5x

x= 800

Bakary 2x= 100+2x

x=100

c. at what price(s) will bread be sold in the two bakeries?

Bakery 1= 800"\times" 4500= 3,600,000

Bakery 2= 100"\times" 4500= 450,000